OpenStudy (ronaldo7):

Find the zero(s) for y= 4x+2 / x-6 Help!! Please

OpenStudy (amking):

Hey. Are you using a graphing calculator for this or by hand?

OpenStudy (ronaldo7):

hello

OpenStudy (shuhbamprakash):

Tell me for what x, y is not defined?

OpenStudy (shuhbamprakash):

To know the zeroes, put y=0 in the equation and solve for x

OpenStudy (photon336):

You can find the zeroes by taking the numerator and denominator and setting each = to zero. \[\frac{ 4x+2 }{ x-6 }\] \[4x+2 = 0 ; x-6 = 0 \]

OpenStudy (solomonzelman):

If you set a denominator=0, then you are really finding the singular points. (When you set x-6=0, you are finding an asymptotic discontinuity.)

OpenStudy (solomonzelman):

and yes if you want to find the zeros of the function, you set numerator=0.